MKTEXMF(1) MKTEXMF(1) NAME mktexmf−createaMetafontsourcefile SYNOPSIS mktexmffont DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualKpathsea:Alibraryforpathsearching. mktexmfisusedtogeneratetheMETAFONTsourcefileforfont,ifpossible.Forexample, ecrm1200orcmr11. Thenameofthegeneratedfileisprintedonstandardoutput. mktexmfistypicallycalledbyotherprograms,ratherthanfromthecommandline. OPTIONS mktexmfacceptsthefollowingoptions: --helpPrinthelpmessageandexit. --version Printversioninformationandexit. SEEALSO mktexpk(1),mktextfm(1). Kpathsea7.32x 4January1998 1